OpenStudy (anonymous):

Simplify. C 17 4 A. 2,380 B. 680 C. 57,120 D. 4,760

OpenStudy (solomonzelman):

What are you simplifying?

OpenStudy (anonymous):

C 17 4

OpenStudy (anonymous):

I dont know how to do this at all

OpenStudy (anonymous):

mohammad.arqum can you help me

OpenStudy (solomonzelman):

You realize that they way it is, it's amorphous?

OpenStudy (anonymous):

??????????????????

OpenStudy (zarkon):

\[_{17}C_{4}=\frac{17!}{4!(17-4)!}\]

OpenStudy (mathstudent55):

\(_nC_r = \dfrac{n!}{r!(n - r)!} \) \(_{17}C_4 = \dfrac{17!}{4!(17 - 4)!} = \dfrac{17!}{4!(17 - 4)! }= \dfrac{17!}{4! 13!} = \) \(= \dfrac{17 \times 16 \times 15 \times 14 \times \cancel{13!}}{4 \times 3 \times 2 \times \cancel{13!} } \)